OpenStudy (anonymous):

What is the simplest form of the expression? ∛108 + ∛1372 - ∛500 A. 4.2∛4 B. 15∛4 C. 5∛4 D. 16∛4

OpenStudy (anonymous):

I think its A ... not sure though

OpenStudy (anonymous):

\[\sqrt[3]{108}=3\sqrt[3]{4}\] \[\sqrt[3]{1372}=7\sqrt[3]{4}\] \[\sqrt[3]{500}=5\sqrt[3]{4}\] 3+7+5=15...

OpenStudy (anonymous):

Thank you :3